1/3 DISTANCE COVERED WITH 10 KMPH , SECOND 1/3 COVERED WITH 20 KMPH AND THE REST 1/3 CPVERED WITH 60 KMPH .HAT IS AVERGE SPEED. 


 


ACC. TO ME 30 KMPH BUT IN SCRA EXAM PAPER IT IS 18 KMPH


Grade:

1 Answers

Nitish Jangid
11 years ago

let Distance is = x Km

then time to cover distance with 10 KmPH = t1 = Distance/Speed = 1/3x / 10 = 1/30x

       time to cover distance with 20 KmPH = t2 = Distance/Speed = 1/3x / 20 = 1/60x

       time to cover distance with 60 KmPH = t3 = Distance/Speed = 1/3x / 60 = 1/ 180x

Total time to cover distance = 1/30x + 1/60x + 1/180x = 10/180x = 1/18x  

Averge Speed = Total distance / Total time to cover the distance

Averge Speed = x / (1/18x) = 18 KMPH